OpenStudy (anonymous):

The projected worth (in millions of dollars) of a large company is modeled by the equation y = 216(1.03)^x. The variable x represent the number of years since 1997. What is the projected annual percent of growth, and what should the company be worth in 2007?

OpenStudy (theopenstudyowl):

This is very easy..... You take the function and plug in the values for x: y=216(1.03)^1997 and y=216(1.03)^2007. After you calculate each value, simply calculate the percent increase from the first value in contrast to the second value! Good luck... Happy Studying!